Please help me with the following question:
(3 + x) / 7 = (3 + x) / 12
Thank you!

Answer by TimothyLamb(4379)

---
(3 + x)/7 = (3 + x)/12
12(3 + x) = 7(3 + x)
36 + 12x = 21 + 7x
5x = -15
x = -3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php